\begin{thebibliography}{99}

\bibitem{Abaecherli_torus}
Angelo Ab{\"a}cherli, \emph{Local picture and level-set percolation of the {Gaussian} free field on a large discrete torus}, Stochastic Processes Appl. \textbf{129} (2019), no.~9, 3527--3546.

\bibitem{AbaechCerny_trees}
Angelo Ab{\"a}cherli and Ji{\v{r}}{\'{\i}} {\v{C}}ern{\'y}, \emph{Level-set percolation of the {Gaussian} free field on regular graphs. {I}. {Regular} trees}, Electron. J. Probab. \textbf{25} (2020), 1--24, Paper no. 65.

\bibitem{AbacherliCerny2}
Angelo Ab{\"a}cherli and Ji{\v{r}}{\'{\i}} {\v{C}}ern{\'y}, \emph{Level-set percolation of the {Gaussian} free field on regular graphs {II}: finite expanders}, Electron. J. Probab. \textbf{25} (2020), 1--39, Paper no. 130.

\bibitem{Bauer19}
Roland Bauerschmidt, Jiaoyang Huang, and Horng-Tzer Yau, \emph{Local {K}esten-{M}c{K}ay law for random regular graphs}, Comm. Math. Phys. \textbf{369} (2019), no.~2, 523--636.

\bibitem{berestycki2018extremes}
Julien Berestycki, {\'E}ric Brunet, Aser Cortines, and Bastien Mallein, \emph{Extremes of branching Ornstein--Uhlenbeck processes}, HAL \textbf{2018} (2018).

\bibitem{BiskupLouidor18}
Marek Biskup and Oren Louidor, \emph{Full extremal process, cluster law and freezing for the two-dimensional discrete {Gaussian} free field}, Adv. Math. \textbf{330} (2018), 589--687.

\bibitem{bovier2005extreme}
Anton Bovier, \emph{Extreme values of random processes}, Lecture Notes Technische Universit{\"a}t Berlin (2005).

\bibitem{bramson2012tightness}
Maury D.~Bramson and Ofer Zeitouni, \emph{Tightness of the recentered maximum of the two-dimensional discrete {Gaussian} free field}, Commun. Pure Appl. Math. \textbf{65} (2012), no.~1, 1--20.

\bibitem{bramson2016convergence}
Maury D.~Bramson, Jian Ding, and Ofer Zeitouni, \emph{Convergence in law of the maximum of the two-dimensional discrete {Gaussian} free field}, Commun. Pure Appl. Math. \textbf{69} (2016), no.~1, 62--123.

\bibitem{CernyIII}
Ji{\v{r}}{\'{\i}} {\v{C}}ern{\'y}, \emph{Giant component for the supercritical level-set percolation of the {G}aussian free field on regular expander graphs}, Comm. Pure Appl. Math. \textbf{76} (2023), no.~11, 3346--3373.

\bibitem{CerTexWin11}
Ji{\v{r}}{\'{\i}} {\v{C}}ern{\'y}, Augusto Teixeira, and David Windisch, \emph{Giant vacant component left by a random walk in a random \(d\)-regular graph}, Ann. Inst. Henri Poincar{\'e}, Probab. Stat. \textbf{47} (2011), no.~4, 929--968.

\bibitem{chiarini2015}
Alberto Chiarini, Alessandra Cipriani, and Rajat Subhra Hazra, \emph{A note on the extremal process of the supercritical {Gaussian} free field}, Electron. Commun. Probab. \textbf{20} (2015), Paper no.~74, 1--10.

\bibitem{chiarini2016}
Alberto Chiarini, Alessandra Cipriani, and Rajat Subhra Hazra, \emph{Extremes of the supercritical {Gaussian} free field}, ALEA Lat. Am. J. Probab. Math. Stat. \textbf{13} (2016), no.~2, 711--724.

\bibitem{Conchon23}
Guillaume Conchon-Kerjan, \emph{Anatomy of a {Gaussian} giant: supercritical level-sets of the free field on regular graphs}, Electron. J. Probab. \textbf{28} (2023), 1--60, Paper no.~35.

\bibitem{DasHazra2019}
Sayan Das and Rajat Subhra Hazra, \emph{Extremal process of the zero-average {Gaussian} free field for \(d\ge3\)}, Stat. Probab. Lett. \textbf{146} (2019), 42--49.

\bibitem{DingRoyZeit17}
Jian Ding, Rishideep Roy, and Ofer Zeitouni, \emph{Convergence of the centered maximum of log-correlated {Gaussian} fields}, Ann. Probab. \textbf{45} (2017), no.~6A, 3886--3928.

\bibitem{DrewitzGalloPrev}
Alexander Drewitz, Gioele Gallo, and Alexis Pr\'evost, \emph{Generating {G}alton-{W}atson trees using random walks and percolation for the {G}aussian free field}, Ann. Appl. Probab. \textbf{34} (2024), no.~3, 2844--2884.

\bibitem{Saloff97}
Laurent Saloff-Coste, \emph{Lectures on finite {Markov} chains}, in: Lectures on probability theory and statistics (Saint-Flour 1996), Springer, 1997, pp.~301--413.

\bibitem{sznitman2015disconnection}
Alain-Sol Sznitman, \emph{Disconnection and level-set percolation for the {Gaussian} free field}, J. Math. Soc. Japan \textbf{67} (2015), no.~4, 1801--1843.

\bibitem{sznitman2016coupling}
Alain-Sol Sznitman, \emph{Coupling and an application to level-set percolation of the {Gaussian} free field}, Electron. J. Probab. \textbf{21} (2016), 1--26, Paper no. 35.

\bibitem{vdH1}
Remco van~der Hofstad, \emph{Random graphs and complex networks. {Volume 1}}, Camb. Ser. Stat. Probab. Math., vol.~43, Cambridge University Press, 2017.

\begin{comment}
    

\bibitem{Abaecherli_torus}
Angelo Ab{\"a}cherli, \emph{Local picture and level-set percolation of the {Gaussian} free field on a large discrete torus}, Stochastic Processes Appl. \textbf{129} (2019), no.~9, 3527--3546.

\bibitem{AbaechCerny_trees}
Angelo Ab{\"a}cherli and Ji{\v{r}}{\'{\i}} {\v{C}}ern{\'y}, \emph{Level-set percolation of the {Gaussian} free field on regular graphs. {I}. {Regular} trees}, Electron. J. Probab. \textbf{25} (2020), 1--24, Paper no. 65.

\bibitem{AbacherliCerny2}
Angelo Ab{\"a}cherli and Ji{\v{r}}{\'{\i}} {\v{C}}ern{\'y}, \emph{Level-set percolation of the {Gaussian} free field on regular graphs {II}: finite expanders}, Electron. J. Probab. \textbf{25} (2020), 1--39, Paper no. 130.

\bibitem{Bauer19}
Roland Bauerschmidt, Jiaoyang Huang, and Horng-Tzer Yau, \emph{Local {K}esten-{M}c{K}ay law for random regular graphs}, Comm. Math. Phys. \textbf{369} (2019), no.~2, 523--636. %\MR{3962004}

\bibitem{berestycki2018extremes}
Julien Berestycki, {\'E}ric Brunet, Aser Cortines, and Bastien Mallein, \emph{Extremes of branching ornstein-uhlenbeck processes}, HAL \textbf{2018} (2018).

\bibitem{bovier2005extreme}
Anton Bovier, \emph{Extreme values of random processes}, Lecture Notes Technische Universit{\"a}t Berlin (2005).

\bibitem{CernyIII}
Ji{\v{r}}{\'{\i}} {\v{C}}ern{\'y}, \emph{Giant component for the supercritical level-set percolation of the {G}aussian free field on regular expander graphs}, Comm. Pure Appl. Math. \textbf{76} (2023), no.~11, 3346--3373. %\MR{4642820}

\bibitem{CerTexWin11}
Ji{\v{r}}{\'{\i}} {\v{C}}ern{\'y}, Augusto Teixeira, and David Windisch, \emph{Giant vacant component left by a random walk in a random {{\(d\)}}-regular graph}, Ann. Inst. Henri Poincar{\'e}, Probab. Stat. \textbf{47} (2011), no.~4, 929--968.

\bibitem{Conchon23}
Guillaume Conchon-Kerjan, \emph{Anatomy of a {Gaussian} giant: supercritical level-sets of the free field on regular graphs}, Electron. J. Probab. \textbf{28} (2023), 1--60 (English), Paper no. 35.

\bibitem{DasHazra2019}
Sayan Das and Rajat~Subhra Hazra, \emph{Extremal process of the zero-average {Gaussian} free field for \(d\geq 3\)}, Stat. Probab. Lett. \textbf{146} (2019), 42--49.

\bibitem{rodriguez2013phase}
Pierre-Fran{\c{c}}ois Rodriguez and Alain-Sol Sznitman, \emph{Phase transition and level-set percolation for the gaussian free field}, Communications in Mathematical Physics \textbf{320} (2013), no.~2, 571--601.

\bibitem{Saloff97}
Laurent Saloff-Coste, \emph{Lectures on finite {Markov} chains}, Lectures on probability theory and statistics. Ecole d'\'et\'e de probabilit\'es de Saint-Flour XXVI--1996. Lectures given at the Saint-Flour summer school of probability theory, August 19--September 4, 1996, Berlin: Springer, 1997, pp.~301--413.

\bibitem{sznitman2015disconnection}
Alain-Sol Sznitman, \emph{Disconnection and level-set percolation for the gaussian free field}, Journal of the Mathematical Society of Japan \textbf{67} (2015), no.~4, 1801--1843.

\bibitem{sznitman2016coupling}
Alain-Sol Sznitman, \emph{Coupling and an application to level-set percolation of the gaussian free field},  (2016).

\bibitem{vdH1}
Remco van~der Hofstad, \emph{Random graphs and complex networks. {Volume} 1}, Camb. Ser. Stat. Probab. Math., vol.~43, Cambridge: Cambridge University Press, 2017.

\end{comment}

\end{thebibliography}

\documentclass[preprint,ECP]{ejpecp} 
\usepackage{graphicx} 
\usepackage{amsmath, relsize} 
\usepackage{todonotes} 
\usepackage{amsthm} 
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{array}
\usepackage{url}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amsthm}
\usepackage{bbm}
\usepackage{mathtools}
\usepackage{xcolor}
\usepackage{tikz}
\usepackage{enumerate}
\usepackage{setspace}
\usepackage{comment}
\usepackage[scr=boondox,  
cal=esstix]   
{mathalpha}

\SHORTTITLE{Extremes of the GFF on regular graphs}

\TITLE{Extremes of the zero-average Gaussian Free Field on random regular graphs
    \thanks{The
research of L.H. was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research
Foundation)  in the framework to the collaborative research center  "Multiscale Simulation Methods for Soft-Matter Systems" (TRR 146) under Project No. 233630050 and Project-ID 443891315 within SPP 2265.The research of C.M.\ was partly funded by the German Research Foundation (project number 443916008, priority program SPP2265). A.K.'s research is funded by the Cusanuswerk (Bischöfliche Studienförderung).}
    } 

\AUTHORS{%
Lisa~Hartung \footnote{University of Mainz, Germany.
    \EMAIL{lhartung@uni-mainz.de}}\orcid{0000-0002-3420-551X}
\and
  	Andreas Klippel \footnote{Technische Universit\"at Darmstadt, Germany.
  		\EMAIL{klippel@mathematik.tu-darmstadt.de}}
\and 
  Christian M\"onch \footnote{
  	\EMAIL{cmoench25@gmail.com}}\orcid{0000-0002-6531-6482}
}

\KEYWORDS{extreme value theory, Gaussian free field, Gumbel limit, random regular graph} 

\AMSSUBJ{60G70} 
\AMSSUBJSECONDARY{60K35; 05C80} 

\SUBMITTED{??????} 
\ACCEPTED{December 13, 2014} 

\VOLUME{0}
\YEAR{2023}
\PAPERNUM{0}
\DOI{10.1214/YY-TN}

\ABSTRACT{We study the extreme value statistics of the zero-average Gaussian free field (GFF) on random $r$-regular graphs and the Gaussian free field on $r$-regular trees. For random $r$-regular graphs of diverging size, for every fixed $r\ge3$, we show that the rescaled extremal point process of the field is asymptotically distributed, in the annealed sense, as a Poisson point process on the line with intensity $e^{-x}\,\mathrm{d}x$. The same limit behaviour is obeyed by the restriction of the GFF on $r$-regular trees to finite subsets of vertices. Our approach relies on a direct Gaussian comparison argument and precise Green function estimates.
	}

\newtheorem{prop}[theorem]{Proposition}

\renewcommand{\P}{\mathbb{P}}

\newcommand{\1}{\mathbbm{1}}
\newcommand{\G}{\mathscr{G}}

\newcommand{\E}{\mathbb{E}}

\newcommand{\R}{\mathbb{R}}

\newcommand{\T}{\mathscr{T}}

\newcommand{\M}{\mathscr{M}}

\newcommand{\N}{\mathbb{N}}

\newcommand{\zigzag}[1]{%
	\begin{tikzpicture}[#1]%
		\draw (0,0) -- (0.5ex,0ex);%
		\draw (0.5ex,0ex) -- (0.5ex,1ex);%
		\draw (0.5ex,1ex) -- (1ex,1ex);%
	\end{tikzpicture}%
}

\newcommand{\Clbrc}{C^\infty_{\mbox{\zigzag{}}}}

\newcommand{\CM}[1]{\todo[color=blue!50!white, inline]{\color{black} \textbf{CM:} #1}}

\numberwithin{equation}{section}

\begin{document}
\raggedbottom


\section{Introduction and main result}
In recent years, properties of level set percolation and local properties of the zero-average Gaussian free field (GFF) on finite graphs and trees have been studied intensively; see, for example, \cite{AbaechCerny_trees,AbacherliCerny2, CernyIII,DrewitzGalloPrev,sznitman2015disconnection,sznitman2016coupling} In the present article, we focus on the extreme value statistics of such fields  on random $r$-regular graphs and identify the proper scaling of the extremes. The only previous results in this direction for regular graphs are the rough upper bound in \cite{Conchon23}, and the work \cite{DasHazra2019} in the rigid setting of the Euclidean torus. Note that whereas \cite{DasHazra2019} relies on a coupling introduced in \cite{Abaecherli_torus}, we develop a different approach to obtain our main result: we combine a general Gaussian comparison principle for point processes (Theorem~\ref{thm:gauss-comp-ppp}) with estimates for the Green's function on random $r$-regular graphs (Section~4.1) that complement the estimates in \cite{AbacherliCerny2}. We believe both of these contributions to be of independent interest.

Throughout, we let $(\G_N,o_N)$ denote a uniformly rooted $r$-regular graph on $N\in 2\N$ vertices. The corresponding distribution on finite graphs is denoted by $\mathbf{P}$ and the graph distances in $\G_N$ are written as $d_{\G_N}(\cdot,\cdot)$. We work under the law
\[
  \P_N \;=\; \int \P^{\G_N}\, \mathrm{d}\mathbf P(\G_N),
\]
where $\P^{\G_N}$ is the process law conditional on the graph realisation $\G_N$. Given $\G_N$, we consider the canonical simple random walk $(X_k)_{k\ge0}$, 
its continuous-time counterpart $(\overline X_t)_{t\ge0}$, 
and the corresponding zero-average GFF $\Psi_N$ on $\G_N$. 
The latter is defined as the centred Gaussian field with covariances given by
\[
  G_{\G_N}(x, y) 
  := \int_0^\infty \big( \P_x^{\G_N}(\overline{X}_t = y) - \tfrac{1}{N} \big)\, \mathrm{d}t, 
  \qquad x, y \in \G_N,
\]
that is, the \emph{zero-average Green function} of $\G_N$. We further set
\begin{equation}\label{eq:rescaling}
  \sigma_r^2 := \frac{r-1}{r-2}, 
  \qquad
  a_N := \sqrt{2\log N} - \frac{\log\log N + \log(4\pi)}{2\sqrt{2\log N}}, 
  \qquad
  b_N := a_N^{-1},
\end{equation}
and define the rescaled field $Z_x := \frac{\Psi_N(x)}{\sigma_r}$, for  $x\in V_N$. The associated extremal point process and centred maximum are given by
\[
  \mathcal P_N := \sum_{x\in V_N}\delta_{\frac{Z_x-a_N}{b_N}},
  \qquad
  M_N := \max_{x\in V_N}\frac{Z_x-a_N}{b_N}.
\]
Let further $\mathcal P_\infty$ be a Poisson point process on $\R$ with intensity $e^{-x}\,\mathrm{d}x$, 
and set
\[
  M_\infty := \max \mathcal P_\infty.
\]
Denoting by $\Rightarrow$ convergence in distribution, our main result can now be stated as follows:
\begin{theorem}[Extremal point process and maximum on the random $r$-regular graph]\label{thm:graph-ext-proc}
Let $r\ge3$. Then, with the notation above, as $N\to\infty$, there exist sets of graphs 
$\Omega_N,\,N\in\N$, such that $\mathbf{P}(\Omega_N^\mathsf{c})=o(1)$ and on $\Omega_N$
\[
  (\mathcal P_N,\,M_N)\ \overset{\P^{\G_N}}\Longrightarrow\ (\mathcal P_\infty,\,M_\infty).
\]
Consequently, $(\mathcal P_N,\,M_N)\ \overset{\P_N}\Longrightarrow\ (\mathcal P_\infty,\,M_\infty)$, 
which in particular implies that $M_N$ is asymptotically Gumbel-distributed, so that
\[
  \P_N\Big( \max_{x\in \G_N} \Psi_N(x)\ \le\ 
  \sigma_r\Big(a_N+t b_N\Big)\Big)
  \ \longrightarrow\ \exp(-e^{-t}).
\]
\end{theorem}

Our proof of Theorem~\ref{thm:graph-ext-proc} relies on two key ingredients.
\begin{itemize}
    \item[(1)] A {\emph Gaussian comparison principle} for the joint convergence of the rescaled extremal process and the maximum, stated and proven in Section 2. In Section 3 we showcase the usefulness of this method by deploying it for the rescaled extremal process and the rescaled maximum  of the Gaussian free field on finite subsets of the $r$-regular infinite tree, see Theorem \ref{thm:tree-ext-proc}.
    \item[(2)] Precise `almost uniform' {\emph Green function estimates}: In $\mathbf{P}$- probability the Green’s function on $\G_N$ can locally be well approximated by the one on the $r-1$-ary tree for asymptotically almost all vertices. See Section 4.1 for the details. 
\end{itemize}
Finally, Section 4.2 contains the proof of Theorem \ref{thm:graph-ext-proc}.

\section{A Gaussian comparison principle}
Our first result is a Gaussian comparison principle for the joint convergence of the extremal process and the maximum, which is instrumental in deriving our main results. Gaussian comparison techniques have proven to be a powerful technique in the study of extremes of Gaussian fields on graphs, cf.\ \cite{berestycki2018extremes,bramson2012tightness,bramson2016convergence,chiarini2015,chiarini2016,DingRoyZeit17,BiskupLouidor18}.
Throughout this section, we let $$X(0)=\big(X_{ij}(0): 1\leq i\leq j, j\in \N\big) \text{ and } X(1)=\big(X_{ij}(1): 1\leq i\leq j, j\in \N\big)$$
denote generic Gaussian triangular arrays which are independent of each other. Their interpolation is given as 
\[
X(h)=\sqrt{1-h}\, X(0)+\sqrt{h}\,X(1),\quad h\in(0,1),
\]
where scalar multiplication and addition are to be understood entry-wise. For given $N\in\N$, we further write $ X_N(h)=\big(X_{1N}(h),\dots, X_{NN}(h)\big)$ for the $N$-th row vector of $X(h)$, assume that each $X_N(h)$ is a centred Gaussian vector in $\R^N$, and denote by 
\[
\Sigma(N,h)=\operatorname{Cov}\big(X_N(h)\big)=(1-h)\operatorname{Cov}\big(X_N(0)\big)+h\operatorname{Cov}\big(X_N(1)\big), \quad h\in [0,1],
\]
the corresponding covariance matrices. 

\begin{theorem}\label{thm:gauss-comp-ppp}
	Assume that, as $N\to\infty,$
	\begin{itemize}
		\item there are sequences $(a_N)_{N\in\N}$ and $(b_N)_{N\in\N}$ such that
		\begin{equation}\label{eq.ppp}
			\left(\sum_{i=1}^{N} \delta_{(X_{iN}(0)-a_N)/b_N}, \max_{1\le i\le N} \frac{X_{iN}(0)-a_N}{b_N}\right)
			\;\Longrightarrow\; \left(\mathcal{P}_\infty, \max \mathcal{P}_\infty\right),
		\end{equation}
		where $\mathcal{P}_\infty$ is a point process on $\R$ with $\mathcal{P}_\infty\big((0,\infty)\big)<\infty$;
		\item for every bounded and open set $S\subset\R$,
		\begin{equation}\label{eq:tozero}
			\frac{1}{b_N^{2}}\int_0^1\!\sum_{\substack{1\le i\ne j\le N}}
			\big|\Sigma_{ij}(N,1)-\Sigma_{ij}(N,0)\big|\,
			\P\left(\tfrac{X_{iN}(h)-a_N}{b_N}\in S,\ \tfrac{X_{jN}(h)-a_N}{b_N}\in S\right)\,\mathrm{d}h
			\ \to\ 0.
		\end{equation}
	\end{itemize}
	Then, as $N\to\infty,$
	\begin{equation}
		\left(\sum_{i=1}^{N} \delta_{(X_{iN}(1)-a_N)/b_N}, \max_{1\le i\le N} \frac{X_{iN}(1)-a_N}{b_N}\right)
		\;\Longrightarrow\; \left(\mathcal{P}_\infty, \max \mathcal{P}_\infty\right).
	\end{equation}
\end{theorem}
Before addressing the proof, we collect a few auxiliary results. We start with a consequence of the Gaussian integration-by-parts formula, see e.g.\ \cite[Lemma 4.1]{bovier2005extreme} for a detailed derivation.
\begin{lemma}\label{lem:gauss-interp}
	Let $Y(0)=\big(Y_1(0),\dots,Y_n(0)\big)$ and $Y(1)=\big(Y_1(1),\dots,Y_n(1)\big)$ be independent centred Gaussian vectors in $\mathbb{R}^n$ with covariance matrices $\Lambda(i)=\big(\Lambda_{ij}(i)\big), i\in{0,1}$ satisfying $\Lambda_{ii}(j)=1$ for all $(j,i)\in\{0,1\}\times\{1,\dots,n\}$. Set further, for $h\in(0,1)$
	\[
	Y(h)= \sqrt{h} Y(1)+\sqrt{1-h}\,Y(0),
	\qquad
	\Lambda(h)= \operatorname{Cov}\big(Y(h)\big),
	\]
	and let $F\in C^{2}(\R^n,\R)$ be any function satisfying the moderate growth condition
	\[
	|F(x)|+\|\nabla F(x)\|+\|\!\operatorname{Hess}F(x)\|\le C\,\big(1+\|x\|\big)^{m},\quad x\in\R^n,
	\]
	for some $C<\infty, m\in\N$, where $\|\cdot\|$ denotes, for definiteness, the $\max$-norm. 
	Then the map $h\mapsto \E\big[F\big(Y(h)\big)\big]$ is continuously differentiable on $[0,1]$ and
	\begin{equation}\label{eq:interp-integral}
		\E\big[F\big(Y(1)\big)-F\big(Y(0)\big)\big]
		\;=\;\frac12\sum_{i,j=1}^{n}\int_{0}^{1}\bigl(\Lambda_{ij}(1)-\Lambda_{ij}(0)\bigr)\,
		\E\big[\partial_{ij}F\big(Y(h)\big)\big]\,\mathrm{d}h,
	\end{equation}
	where $\partial_{ij}F, 1\leq i,j\leq n$, denote the entries of $\operatorname{Hess} F$.
\end{lemma}

	   
We say that a function $\varphi:\R\to\R$ has \emph{l-bounded support} if there exists $K<\infty$, such that $\varphi(x)=0$ for all $x\leq-K$. A differentiable function $\varphi$ is \emph{eventually constant (on the right)} if $x\mapsto\varphi'(-x)$ has l-bounded support. We denote by $\Clbrc$ the cone of all non-decreasing, eventually constant, smooth functions with l-bounded support. Note that $\varphi\in\Clbrc$ implies that $\operatorname{supp}(\varphi')$ is a bounded open set and, conversely, for every bounded open set $S$ we can find a function $\varphi\in \Clbrc$ such that $\operatorname{supp}(\varphi')=S$. 

\begin{lemma}[{\cite[Lemma 4.1, parts (i),(ii), and (iv)]{berestycki2018extremes}}]\label{lem:pp-equivalence}
	
	Let $(P_t: t\in \N\cup\{\infty\})$ denote a family of point processes on $\mathbb{R}$ where $P_\infty\big((0,\infty)\big) < \infty$ almost surely.  
	The following statements are equivalent:
	\begin{enumerate}[(i)]
		\item $(P_t, \max P_t) \Longrightarrow(P_\infty, \max P_\infty)$ jointly, as $t\to\infty$;
		\item $P_t \Longrightarrow P_\infty$ and $\max P_t \Longrightarrow \max P_\infty$, as $t\to\infty$;
		\item $\mathbb{E}\left[e^{-\langle P_t,\varphi \rangle}\right] \;\to\; 
		\mathbb{E}\left[e^{\langle P_\infty,\varphi\rangle}\right],$ as $t\to\infty$, 
		for all $\varphi\in\Clbrc$.
	\end{enumerate}
\end{lemma}

\begin{proof}[Proof of Theorem~\ref{thm:gauss-comp-ppp}]
By Lemma~\ref{lem:pp-equivalence} and~\eqref{eq.ppp}, it suffices to prove that
\[
\Big|
\E e^{-\langle \mathcal{P}_{N}(1),\varphi\rangle}
-
\E e^{-\langle \mathcal{P}_{N}(0),\varphi\rangle}
\Big|
\longrightarrow 0
\quad\text{for all }\varphi\in\Clbrc.
\]
Fix $\varphi\in\Clbrc$ and set $S=\operatorname{supp}(\varphi')$.
For $x=(x_1,\dots,x_N)\in\R^N$, define
\[
F_N(x):=\exp\Big(-\sum_{k=1}^N \varphi\Big(\frac{x_k-a_N}{b_N}\Big)\Big),
\quad\text{so that}\quad
e^{-\langle \mathcal{P}_N(h),\varphi\rangle}=F_N(X_N(h)).
\]

Let $\sigma_{iN}^2=\operatorname{Var}(X_{iN}(0))=\operatorname{Var}(X_{iN}(1))$ and put
$Z_{iN}(h)=X_{iN}(h)/\sigma_{iN}$.  
Then $Z_N(0)$ and $Z_N(1)$ are independent centred Gaussian vectors
with unit variance, hence all $Z_N(h)=\sqrt{1-h}\,Z_N(0)+\sqrt{h}\,Z_N(1)$ have unit variance. Defining
\[
\widetilde F_N(z)
:=F_N(\sigma_{1N}z_1,\dots,\sigma_{NN}z_N)
=\exp\Big(-\sum_{k=1}^N 
\varphi\Big(\frac{\sigma_{kN}z_k-a_N}{b_N}\Big)\Big),
\]
we have $F_N(X_N(h))=\widetilde F_N(Z_N(h))$.  
Applying Lemma~\ref{lem:gauss-interp} to $(Z_N(h),\widetilde F_N)$ gives
\[
\E[\widetilde F_N(Z_N(1))-\widetilde F_N(Z_N(0))]
=\frac12\sum_{i\ne j}(\tilde\Lambda_{ij}(N,1)-\tilde\Lambda_{ij}(N,0))
\int_0^1\E[\partial_{ij}^2\widetilde F_N(Z_N(h))]\,\mathrm{d}h,
\]
with $\tilde\Lambda_{ij}(N,h)=\Sigma_{ij}(N,h)/(\sigma_{iN}\sigma_{jN})$.
Since for $i\ne j$
\[
\partial_{ij}^2\widetilde F_N(z)
=\frac{\sigma_{iN}\sigma_{jN}}{b_N^2}\,
\varphi'\Big(\frac{\sigma_{iN}z_i-a_N}{b_N}\Big)
\varphi'\Big(\frac{\sigma_{jN}z_j-a_N}{b_N}\Big)
\widetilde F_N(z),
\]
the factors $\sigma_{iN}\sigma_{jN}$ cancel, and

\begin{align*}
&\Big|\E[F_N(X_N(1))-F_N(X_N(0))]\Big|\\
&\le\frac{\|\varphi'\|_\infty^2}{2b_N^2}
\int_0^1\sum_{i\ne j}
|\Sigma_{ij}(N,1)-\Sigma_{ij}(N,0)|
\P\Big(\tfrac{X_{iN}(h)-a_N}{b_N}\in S,\tfrac{X_{jN}(h)-a_N}{b_N}\in S\Big)\,\mathrm{d}h.
\end{align*}
By~\eqref{eq:tozero}, the right-hand side tends to $0$ as $N\to\infty$,
yielding the claim.
\end{proof}

The following simple analytical lemma provides an upper bound for the probability in \eqref{eq:tozero}, which we use to apply Theorem~\ref{thm:gauss-comp-ppp}.
\begin{lemma}\label{lem:doublegauss}
	Let $(X,Y)$ be a centered bivariate Gaussian vector with unit variances and correlation $\rho \in [0,1)$. 
	Let $S \subset \mathbb{R}$ be a bounded open set with Lebesgue measure $|S| < \infty$. 
    For $u > 0$, define
	\[
	d(u,S) := \inf_{s \in S} |u + s| = \operatorname{dist}(-u, S).
	\]
	Then
	\[
	\mathbb{P}\big(X - u \in S,\; Y - u \in S\big)
	\;\le\;
	\frac{|S|^{2}}{2\pi\sqrt{1-\rho^{2}}}\,
	\exp\!\Big(-\frac{d(u,S)^{2}}{1+\rho}\Big).
	\]
\end{lemma}

\begin{proof}
We have
	\[
	\mathbb{P}(X-u\in S,\ Y-u\in S)
	= \iint_{(u+S)\times(u+S)} f(x,y)\,dx\,dy,
	\]	
    where 
	\[
	f(x,y)
	= \frac{1}{2\pi\sqrt{1-\rho^{2}}}
	\exp\!\left(-\frac{x^{2}-2\rho xy+y^{2}}{2(1-\rho^{2})}\right),
	\]
	ist the joint density of $X$ and $Y$. For $\rho \in [0,1)$, we have the elementary inequality
	\[
	x^{2}-2\rho xy+y^{2} \ge (1-\rho)(x^{2}+y^{2}),
	\]
	since the difference equals $\rho(x-y)^{2} \ge 0$. 
	For every $x \in u+S$, we have $|x| \ge d:=d(u,S)$, and the same holds for $y \in u+S$. Therefore,
\begin{equation}\label{eq:inequality}
	\inf_{(x,y)\in (u+S)\times(u+S)} 
	\frac{x^{2}-2\rho xy+y^{2}}{2(1-\rho^{2})}
	\ge
	\frac{d^{2}+d^{2}}{2(1+\rho)}
	= \frac{d^{2}}{1+\rho}.
\end{equation}
	Using \eqref{eq:inequality} to uniformly bound the density and oberving that the integration region has area $|S|^{2}$, the claim follows.
\end{proof}
\section{Extrema of the Gaussian Free field on the regular tree}
Our first result about the extremal behaviour of the GFF concerns the infinite $r$-regular tree $\T=\T_r$ for $r\geq 3$. Let $X=(X_k)_{k=0}^\infty$ denote simple random walk on $\T$ and let $\P_x$ and $\E_x$ denote its law and expectation given $X_0=x$. The asscociated Green function is then given by
\begin{equation}
	g(x,y) := {\E}_x\left[\sum_{k=0}^\infty \mathbbm{1}_{\{X_k = y\}}\right] = \frac{r-1}{r-2} \left(\frac{1}{r-1}\right)^{d_{\T}(x,y)} \quad \text{for } x, y \in \T,
\end{equation}
where $d_{\T}(\cdot,\cdot)$ denotes graph distance in $\T$. Let $\psi=\big(\psi(x), x\in \T\big)$ denote the GFF on $\T$, i.e.\ the centred Gaussian field with covariance
\[
\E[\psi(x)\psi(y)]=g(x,y),\quad x,y\in \T
\]
Throughout, we assume that $V_N\subset\T$ is a subtree with $N$ vertices.  Recall that $\sigma_r^2$
, $a_N$ and $b_N$ are defined in \eqref{eq:rescaling}.
Define the centred and rescaled field
\[
Z_x \;:=\; \frac{\psi(x)}{\sigma_r},
\qquad x\in V_N,
\]
so that $\operatorname{Var}(Z_x)=1$.  
Let
\[
\mathcal P_N \;:=\; \sum_{x\in V_N}\delta_{\frac{Z_x-a_N}{b_N}},
\qquad 
M_N \;:=\; \max_{x\in V_N}\frac{Z_x-a_N}{b_N}.
\]
Recall that  $\mathcal P_\infty$ is a Poisson point process on $\R$ with intensity $e^{-x}\,\mathrm{d}x$, and
$
M_\infty:=\max \mathcal P_\infty.
$

\begin{theorem}[Extremal point process and maximum on the $r$-regular tree]\label{thm:tree-ext-proc}
	Let $r\ge3$, 
	then, as $N\to\infty$,
	\[
	(\mathcal P_N,\,M_N)\ \Longrightarrow\ (\mathcal P_\infty,\,M_\infty).
	\]
	In particular, $M_N$ is asymptotically standard-Gumbel distributed.
\end{theorem}

\begin{proof}
	We wish to apply Theorem~\ref{thm:gauss-comp-ppp}.  
	Recall that
	\[
	\rho_{xy}:=\E[Z_xZ_y]
	=\left(\frac{1}{r-1}\right)^{d_\T(x,y)}\in[0,1)
	\qquad \text{ for all }x,y\in V_N \text{ with }x\neq y.
	\]
Note that $\rho_1 = 1/(r-1) < 1$ if and only if $r \ge 3$. 
	Let $(X_x(0))_{x\in V_N}$ be i.i.d.\ standard Gaussian variables, independent of $\psi$, and define
$	
	X_x(1):=Z_x,\; x\in V_N.
$
	For $h\in[0,1]$ set
	\[
	X_x(h):=\sqrt{1-h}\,X_x(0)+\sqrt{h}\,X_x(1),\qquad
	X_N(h):=(X_x(h))_{x\in V_N}\in\R^N .
	\]
	Let $(a_N,b_N)$ be as in \eqref{eq:rescaling}.
By classical extreme value theory, see e.g.\ \cite[Thm.\ 5.5.2]{bovier2005extreme}, we have
\[
\Bigg(\sum_{x\in V_N} \delta_{(X_x(0)-a_N)/b_N},\ 
\max_{x\in V_N}\frac{X_x(0)-a_N}{b_N}\Bigg)
\Longrightarrow (\mathcal P_\infty,\max \mathcal P_\infty),
\]
and consequently condition~\eqref{eq.ppp} is satisfied.
It remains to show that \eqref{eq:tozero} holds, i.e.\ that for any fixed but arbitrary bounded open set $S\subset\R$ with $|S|<\infty$
\[
T_N
:= 
\frac{1}{b_N^{2}}\int_0^1\sum_{x\ne y}
\big|\Sigma_{xy}(N,1)-\Sigma_{xy}(N,0)\big|\,
\P\Big(\tfrac{X_{x}(h)-a_N}{b_N}\in S,\,
\tfrac{X_{y}(h)-a_N}{b_N}\in S\Big)\, \mathrm{d}h
\]
vanishes as $N\to\infty$. To this end, we first observe that, for $x\ne y$ and $h\in[0,1]$, 
$(X_{x}(h),X_{y}(h))$ is a centred Gaussian vector with unit variances
and correlation $\rho_{xy}(h)=h\rho_{xy}\in[0,1)$. We next apply Lemma~\ref{lem:doublegauss}, first noting that $(X-u)/b \in S$ is equivalent to $X \in u + bS$, and that 
	$|bS| = b|S|$ while $\operatorname{dist}(-u,bS) = b\,\operatorname{dist}(-u/b,S)$, to obtain
	\begin{equation*}
		\P\Big(\tfrac{X_{x}(h)-a_N}{b_N}\in S,\,
		\tfrac{X_{y}(h)-a_N}{b_N}\in S\Big)
		\ \le\ 
		\frac{b_N^{2}|S|^{2}}{2\pi\sqrt{1-\rho_{xy}(h)^2}}\,
		\exp\Big(-\frac{a_N^2}{1+\rho_{xy}(h)}\Big),
	\end{equation*}
	and thus
	\[T_N
	\le 
	\frac{|S|^{2}}{2\pi}\int_0^1\sum_{x\ne y}
	\big|\Sigma_{xy}(N,1)-\Sigma_{xy}(N,0)\big|\, \frac{1}{\sqrt{1-\rho_{xy}(h)^2}}\,
	\exp\Big(-\frac{a_N^2}{1+\rho_{xy}(h)}\Big)\, \mathrm{d}h.
	\]
	
	Observe that $\rho_{xy}(h)=h\rho_{xy}\le\rho_{xy}$ and that $H(\rho):=(1-\rho^2)^{-1/2}e^{-a_N^2/(1+\rho)}$ is strictly increasing, which leads to
	
	\begin{equation} \label{eq:T_Nbound}
		T_N
		\ \le\ 
		\frac{|S|^2}{2\pi}\sum_{x\ne y}\rho_{xy}\,H(\rho_{xy})
		=
		\frac{|S|^2}{2\pi}\sum_{x\ne y}\rho_{xy}\,
		\frac{1}{\sqrt{1-\rho_{xy}^2}}\,
		\exp\Big(-\frac{a_N^2}{1+\rho_{xy}}\Big).
	\end{equation}
	Grouping unordered pairs of vertices by mutual tree distance, we set
	\[
	N_k := \big\{\{x,y\}\subset V_N : d_\T(x,y)=k\big\}
	\quad \text{and}\quad \rho_k=(r-1)^{-k}.
	\]
	Then, using monotonicity of $H(\cdot)$ again,
	\begin{equation}\label{eq:sumbound592}
		\sum_{x\ne y}\rho_{xy}\,H(\rho_{xy}) = \sum_{k\ge1}\sum_{\{x,y\}\in N_k} \rho_{k}\,H(\rho_{k})=\sum_{k\ge1}|N_k|\rho_{k}\,H(\rho_{k})\leq H(\rho_1)\sum_{k\ge1}|N_k|\rho_{k}.
	\end{equation}
	For each $x\in V_N$, the sphere of radius $k$ contains at most $r(r-1)^{k-1}$ vertices, but never more than~$N$,
	which implies
	\[
	\big|\{y\in V_N : d_\T(x,y)=k\}\big| \le \min\big\{r(r-1)^{k-1},\,N\big\}.
	\]
	Summing over $x$ gives
	\begin{equation}\label{eq:N_kbound}
		|N_k| \le C\,N\,\min\{r(r-1)^{k-1},\,N\},
	\end{equation}
	for some constant $C<\infty$. Inserting \eqref{eq:N_kbound} into \eqref{eq:sumbound592}, combining the resulting bound with \eqref{eq:T_Nbound} and using that $\rho_k=(r-1)^{-k}$ yields
	\begin{align*}
		T_N
		& \le \frac{CN|S|^2 H(\rho_{1})}{2\pi}\sum_{k\geq 1}\rho_k\min\big\{r(r-1)^{k-1},\,N\big\}\\
        & = \frac{CN|S|^2 H(\rho_{1})}{2\pi}\sum_{k\geq 1}\min\big\{r(r-1)^{-1},\,N(r-1)^{-k}\big\}\\
		& = \frac{CN|S|^2 H(\rho_{1})}{2\pi}\left( \sum_{k\leq \log_{r-1}N}\frac r{r-1} + N\sum_{k > \log_{r-1}N}(r-1)^{-k}\right)\\
		& \leq \tilde C\,H(\rho_1)\,  N\log N,
	\end{align*}
	for some $\tilde C<\infty$ depending only on $S$. Using $a_N^2=2\log N+O(\log\log N)$ we get
\[
T_N 
\le N^{\,1-\frac{2}{1+\rho_1}}(\log N)^{O(1)}
= N^{\,\frac{2-r}{r}}(\log N)^{O(1)}.
\]
Since $\frac{2-r}{r}<0$ if $r\ge3$ we have $T_N=o(1)$ and Theorem~\ref{thm:gauss-comp-ppp} proves the claim. 
\end{proof}

\section{Extrema of the Gaussian Free field on the random regular graphs}
\subsection{The zero-average Green function on $\G_N$}
Approximations for the Green function on random $r$-regular graphs are rather well-studied, 
see e.g.\ \cite{Bauer19,AbacherliCerny2,Conchon23}. In this section, we prove that, for most vertices, the the Green function can be computed precisely. We collect a few necessary results in the following proposition, where we use the notation \( B_k(x, \G_N) \) for the $k$-neighbourhood of vertex $x$ in $\G_N$.
\begin{prop}\label{prop:collection}
	There exist $k_1,k_2,k_3>0$, and $K_1<\infty$ such that the following properties are satisfied for the random graphs $\G_N$ with high probability under $\mathbf{P}$ as $N\to\infty$:
	\begin{enumerate}[(i)]
		\item $\G_N$ is a $k_1$-expander, and in particular connected.
		\item $ B_{k_1\log N}(x,\G_N)$ contains at most one cycle for all $x\in[N]$.
		\item The spectral gap of $\G_N$ is at least $k_2.$
		\item For all $x,y\in[N]$ we have $$G_{\G_N}(x,y)\leq K_1 (r-1)^{-d_{\G_N}(x,y)}\vee N^{-k_3}.$$
	\end{enumerate}
\end{prop}
\begin{proof}
	Items (i) and (ii) are taken from \cite[Prop.\ 2.1]{Conchon23}. Item (iii) is discussed in \cite{CerTexWin11} and item (iv) is an immediate consequence of the upper bound given in \cite[Prop.\ 2.2]{AbacherliCerny2}.
\end{proof}

We say that a vertex \( x \in [N] \) is \emph{\( \ell \)-good} if its \( \ell \)-neighbourhood \( B_\ell(x; \G_N) \) in \( \G_N \) is isomorphic (as a rooted graph) to the \( \ell \)-neighbourhood \( B_\ell(o; \T) \) of the root in the infinite $r$-regular tree. That is, \( x \) has no cycles within distance \( \ell \), and therefore the local structure around \( x \) is tree-like up to depth \( \ell \). We further call a vertex \( x \in [N] \) \emph{\( \ell \)-bad} if it is not \( \ell \)-good. Let
\[
\mathrm{bad}_\ell^N := \left\{ x \in [N] : x \text{ is } \ell\text{-bad} \right\}
\]
denote the set of all \( \ell \)-bad vertices in \( \G_N \).
\begin{lemma}\label{lem:badlem}
	Let $\ell=\ell(N)<\log_{r-1}(5N)$. There exists some constant $K_2<\infty$ such that for the set $\mathrm{bad}_\ell^N$ of $\ell$-bad vertices and any $z$ the following is true:
	\begin{equation}\label{eq:badbound}
		\mathbf{P}(|\mathrm{bad}_\ell^N|\geq r^2(r-1)^{2\ell-2} z)\leq \frac{K_2}{z}.    
	\end{equation}
	
\end{lemma}
\begin{proof}
	The proof is similar in vein to that of \cite[Lemma 4.2]{vdH1}, here are the details: The graph $\G_N$ has the same distribution as the \emph{configuration model} $\M_N$ derived from the degree sequence $(r,\dots,r)$ of length $N$ conditioned on being loop-free and simple. $\M_N$ can be constructed in the following way: to every $x\in [N]$ attach $r$ formal objects $h_1(x),\dots,h_r(x)$ called \emph{half-edges}. Now pick a uniform matching on the set $\{h_i(x),1\leq i\leq r, 1\leq x \leq N\}$ and identify a matched pair of half-edges with an edge (note that self-loops are allowed). The resulting collection of edges can be identified with a multigraph $\M_N$ on the vertex set $[N]$. It is well-known, see e.g.\ \cite{vdH1}, that the probability that $\M_N$ is loop-free and simple is non-vashing as $N\to\infty$, implying that results which hold with high probability for $\M_N$ also occur with high probability for $\G_N$. We explore the neighbourhood of vertex $1$ in $\G_N$ sequentially by revealing half-edges during a breadth-first search. More precisely, set $\mathrm{exp}_0=\{1\}$ and $\mathrm{act}_0=\{h_1(1),\dots,h_r(1)\}$ to initialise the exploration. The exploration process\footnote{Note that the matching generating $\M_N$ can also be \emph{constructed} sequentially starting with the half-edges incident to $1$ and the resulting process has the same distribution as the exploration described here.} now updates the explored vertices $\mathrm{exp}_t$ and the active half-edges $\mathrm{act}_t$ iteratively:
	\begin{itemize}
		\item Let $x_{t+1}$ denote the vertex closest to $1$ in $\G_N$ that can be reached by a path using only vertices in $\mathrm{exp}_t$ and has a half-edge in $\mathrm{act}_t$ (if there are several such vertices we use the one with smallest index) and then let $H_{t+1}$ denote this half-edge (again, if there are more than one, we choose the one with the smallest index).
		\item Reveal the half-edge $H'_{t+1}$ that $H_{t+1}$ has been matched to.
		\item Add the endpoint $y_{t+1}$ of $H'_{t+1}$ to $ \exp_{t}$ to get $\exp_{t+1}$, then remove $H_{t+1},H'_{t+1}$ from $\mathrm{act}_t$ and additionally add all unrevealed half-edges of $y_t$ to $\mathrm{act}_t$ to obtain $\mathrm{act}_{t+1}$.
	\end{itemize}

	We say that a \emph{collision} occurs in exploration step $t$, if $H'_t\in \mathrm{act}_t$. Let $\tau_{\mathrm{coll}}$ denote the first exploration step with a collision. The subgraph of $\M_N$ consisting of explored vertices and revealed edges during exploration steps $1,\dots,\tau_{\mathrm{coll}}-1$ is a tree. Using the assumption on the upper bound on $\ell$ and the fact that $|\mathrm{act}_t|\leq r+(r-1)t$ for all $t$, it follows that
	\begin{align*}
		\mathbf{P}(1\in \mathrm{bad}_\ell^N) & = \mathbf{P}(\tau_{\mathrm{coll}}\leq r(r-1)^{\ell-1})\leq \sum_{t=1}^{r(r-1)^{\ell-1}}\frac{r+(r-1)t}{2rN-2t}\\
		&\leq \frac{2}{N}\sum_{t=1}^{r(r-1)^{\ell-1}}{r+(r-1)t}\leq \frac{1}{N}\,K r^2(r-1)^{2\ell-2},
	\end{align*}
	for some $K<\infty.$ By symmetry, we find that $\mathbf{E}[|\mathrm{bad}_\ell^N|]\leq K r^2(r-1)^{2\ell-2}$ and thus \eqref{eq:badbound} follows by Chebyshev's inequality.
\end{proof}

Next, we recall that the continuous time simple random walk $\overline{X}$ on a graph $G$ mixes exponentially fast if $G$ has a positive spectral gap.
\begin{lemma}[{\cite[Corollary\ 2.1.5.]{Saloff97}}]\label{lem:Saloff} Let $G$ be a connected $r$-regular graph of size $N$ with spectral gap $\kappa>0$, then for any vertices $x,y\in G$
	\[
	\Big|\P_x^G(\overline{X}_t=y)-\frac{1}{N}\Big|\leq \textup{e}^{-\kappa\, t} \quad\text{ for all }t>0.
	\]
	
\end{lemma}
Below, we will see that for the proof of Theorem~\ref{thm:graph-ext-proc} we only require precise control on the Green function of the $\ell_N$-good vertices which is our next goal.
\begin{lemma}\label{lem:goodgreenbound}
	Fix $\ell_0\in\N$ and let $\tilde\ell_N=\lfloor \varepsilon\log N \rfloor$ with $\varepsilon>0$. There exists a constant $k_4=k_4(\ell_0,r,\varepsilon)>0$ such that for all vertices $y$ at graph distance at most $\ell_0$ of an $\tilde\ell_N$-good vertex $x$
	\begin{equation}\label{eq:localapprox1}
		\lim_{N\to\infty}\mathbf P\big(|G_{\G_N}(x,y)-g(o,y')|\leq N^{-k_4}\big)=1,
	\end{equation}
	where $y'$ has depth $d_{\G_N}(x,y)$ in $\T$.
\end{lemma}
\begin{proof}
	For $\ell\in\N$, let $g_\ell$ denote the Green function of continuous time simple random walk on $\T$ started at the root and killed upon reaching depth $\ell+1$. Let $\tau_N=\tau_N(x)$ denote the first time that $\overline{X}$ visits $B_{\tilde\ell_N-1}(x;\G_N)^\mathsf{c}$ when started in $x\in\G_N$. Because $x$ is assumed to be $\tilde\ell_N$-good, there is a perfect coupling between $\overline{X}$ under $\P_x^{\G_N}$ and under $\P_o^{\T}$ up until $\tau_N$ and we have
	\begin{align*}
		\big|G_{\G_N}(x,y)-g(o,y') \big| & \leq \Big|\E_x^{\G_N}\Big[\int_0^{\tau_N}\1\{\overline{X}_t=y\}-\frac{1}{N}\,\textup{d}t\Big]-g_{\ell_N-1}(o,y') \Big|\\
		&\phantom{\leq\;}+ \E_x^{\G_N}\Big[\int_{\tau_N}^{\infty}\1\{\overline{X}_t=y\}-\frac{1}{N}\,\textup{d}t\Big] +|g(o,y')-g_{\ell_N-1}(o,y')|\\
		&\leq \frac{\E_x^{\G_N}[\tau_N]}{N} + \E_x^{\G_N}\Big[\int_{\tau_N}^{\infty}\1\{\overline{X}_t=y\}-\frac{1}{N}\,\textup{d}t\Big] +|g(o,y')-g_{\ell_N-1}(o,y')|.
	\end{align*}
	The first term is of order $\log N/N$, since $\overline{X}$ has a linear drift upwards before $\tau_N$, and the third term is clearly decaying exponentially in $\tilde\ell_N-1$. Consequently, it remains to bound the term
	\[
	e_N=\E_x^{\G_N}\Big[\int_{\tau_N}^{\infty}\1\{\overline{X}_t=y\}-\frac{1}{N}\,\textup{d}t \Big].    
	\]
	We have, for suitably chosen small $\delta>0$,
	\begin{align*}
		e_N\leq \frac{1}{k_2}N^{-\delta k_2} +\E_x^{\G_N}\Big[\1\{\tau_N < \delta\log N\}\int_{\tau_N}^{\delta\log N}\1\{\overline{X}_t=y\}-\frac{1}{N}\,\textup{d}t\Big]
	\end{align*}
	with high $\mathbf{P}$-probability by Proposition~\ref{prop:collection} and Lemma~\ref{lem:Saloff}. Additionally,
	\[
	\E_x^{\G_N}\Big[\1\{\tau_N < \delta\log N\}\int_{\tau_N}^{\delta\log N}\1\{\overline{X}_t=y\}-\frac{1}{N}\,\textup{d}t\Big]\leq \delta\log N\; \P(Y(\delta,N)>\ell_N),
	\]
	where $Y(\delta,N)\sim\operatorname{Poisson}(\delta\log N)$ counts the jumps of $\overline{X}$ up to time $\delta\log N$, and thus the remaining bound follows from Poisson concentration, which concludes the proof.
\end{proof}

\subsection{Proof of main result}
    \begin{proof}[Proof of Theorem~\ref{thm:graph-ext-proc}]
     Define $\Omega'_N$ as the event that $\G_N$ satisfies 
    	\ref{prop:collection}\textup{(i)}–\textup{(iv)} and the high probability event in the conclusion of Lemma~\ref{lem:goodgreenbound} applied with $\varepsilon=\varepsilon_0$ and $\ell_0=2$, where
        
    \[
    \varepsilon_0:=\frac{\sigma_r^2}{8 K_1\log(r-1)},\quad\text{and}\quad \ell_N:=\lfloor \varepsilon_0\log N \rfloor.
    \]
        We get by choosing $z=N^\alpha$ with fixed $\alpha\in(0,1)$ in Lemma~\ref{lem:badlem}, that
    	\[
    	\Omega''_N=\big\{|\text{bad}_N^{\ell_N}|\ \le\ N^\alpha (r-1)^{2\ell_N}\big\}\qquad \text{occurs with $\mathbf P$-probability }1-o(1).
    	\]
    	From now on, we work on a realisation $\G_N\in \Omega_N=\Omega'_N\cap\Omega''_N$, such that in particular $|\text{bad}_N^{\ell_N}|=o(N)$. 
    	Since also $\sqrt{\operatorname{Var}(\Psi_N(x))}=\sigma_r(1+O(N^{-k_4}))$, we have for 
    	\(
    	\widehat Z_x  :={\Psi_N(x)}/{\sqrt{\operatorname{Var}(\Psi_N(x))}},
    	\) that
    	\begin{equation}\label{eq:witherror}
    	\Psi_N(x)>\sigma_r\Big(a_N+\tfrac{t}{a_N}\Big)
    	\quad\Longleftrightarrow\quad
    	\widehat Z_x>a_N+\tfrac{t}{a_N}+\delta_N,\quad x\in \G_N,
    	\end{equation}
    	where $\delta_N=o(a_N^{-1})$ is a lower order error term.
    	By choice of $a_N$ and a Gaussian tail bound, we have $\bar\Phi(a_N+t/a_N+o(1))=N^{-1+o(1)}$, and by a union bound ,
    	\[
    	\P^{\G_N}\Big(\max_{x\in \text{bad}_N^{\ell_N}} \Psi_N(x) > u_\Psi\Big)
    	\ \le\ |\text{bad}_N^{\ell_N}|\,\bar\Phi(a_N+t/a_N+o(1))
    	\ =\ o(N)\cdot N^{-1+o(1)}\ =\ o(1),
    	\]
        where  \( u_\Psi = \sigma_r\big(a_N+\tfrac{t}{a_N}\big)\). In particular, it suffices to control the extremes over the $\ell_N$-good vertices, which we will denote by $V_N$. We define the rescaled point processes
    	\[
    	P_{\widehat Z}=\sum_{x\in V_N}\delta_{(\widehat Z_x-a_N)/b_N},\qquad 
    	P_\eta=\sum_{x\in V_N}\delta_{(\eta_x-a_N)/b_N},
    	\]
    	where $(\eta_x)_{x\in V_N}$ are i.i.d.\ $N(0,1)$, independent of $\Psi_N$.
        For $x\neq y\in \G_N$,
    	\[
    	\rho_{xy}\ :=\ \E[\widehat Z_x \widehat Z_y]
    	=\frac{G_{\G_N}(x,y)}{\sqrt{\operatorname{Var}(\Psi_N(x))\,\operatorname{Var}(\Psi_N(y))}}
    	= \frac{G_{\G_N}(x,y)}{\sigma_r^2}\big(1+O(N^{-k_4})\big).
    	\]
    	Using Proposition~\ref{prop:collection}\textup{(iv)} and $\sigma_r^2>0$, there exists $C<\infty$ such that
    		\[
    	0\le \rho_{xy}\ \le\ C\,(r-1)^{-d_{\G_N}(x,y)}\ \vee\ C\,N^{-k_3}.
    	\]
    	Let $S\subset\R$ be a bounded open set. For $h\in[0,1]$, the pair
    \[
    	\big(X_{x}(h),X_{y}(h)\big)
    	=\big(\sqrt{1-h}\,\eta_x+\sqrt{h}\,\widehat Z_x,\ \sqrt{1-h}\,\eta_y+\sqrt{h}\,\widehat Z_y\big)
    \]
    is a centred, unit‐variance Gaussian vector with correlation $\rho_{xy}(h)=h\,\rho_{xy}$. We now argue precisely as in the proof of Theorem~\ref{thm:tree-ext-proc}, aiming to bound
    
    %
    \[
    T_N(\G_N)
    :=
    \frac{1}{b_N^{2}}\int_0^1\sum_{x\ne y}
    \big|\Sigma_{xy}(N,1)-\Sigma_{xy}(N,0)\big|\,
    \P\Big(\tfrac{X_{x}(h)-a_N}{b_N}\in S,\ \tfrac{X_{y}(h)-a_N}{b_N}\in S\Big)\,\mathrm{d}h.
    \]
    Setting $\Sigma_{xy}(N,1)-\Sigma_{xy}(N,0)=\rho_{xy}$, recalling that 
    $q\mapsto H(q)=(1-q^2)^{-1/2}e^{-a_N^2/(1+q)}$ is increasing on $[0,1)$ and that 
    $\rho_{xy}(h)\le\rho_{xy}$, we obtain from an application of Lemma~\ref{lem:doublegauss} that
    \[
    \frac{1}{b_N^{2}}\P\Big(\tfrac{X_{x}(h)-a_N}{b_N}\in S,\ \tfrac{X_{y}(h)-a_N}{b_N}\in S\Big)
    \ \le\
    C_S\,H(\rho_{xy}),
    \qquad C_S:=\tfrac{|S|^2}{2\pi}.
    \]
    Consequently,
    \[
    T_N(\G_N)\ \le\ C_S\sum_{x\ne y}\rho_{xy}\,H(\rho_{xy}).
    \]
    
    	As on the tree, we group unordered pairs by $k=d_{\G_N}(x,y)$ and set
    	\[
    	n_k := \big|\{\{x,y\}\subset V_N : d_{\G_N}(x,y)=k\}\big|, \qquad \rho_k:=\sup\{\rho_{xy}:d_{\G_N}(x,y)=k\},\quad k\geq 1.
    	\]
    	By Proposition~\ref{prop:collection}\textup{(i)}–(ii), there exist $C_1,C_2>0$ such that
    	\begin{equation}\label{eq:nk-rhok-bounds}
    		n_k \le C_1 N (r-1)^k,\qquad 
    		\rho_k \le C_2 (r-1)^{-k},\qquad 1\le k\le \operatorname{diam}(\G_N),
    	\end{equation}
    	and $\operatorname{diam}(\G_N)=O(\log N)$ by Proposition~\ref{prop:collection}\textup{(i),(iii)}.
    	Now fix $\delta>0$ and set
    	\[
    	k_\star \;=\; \Big\lceil \frac{k_3+\delta}{\log(r-1)} \,\log N \Big\rceil.
    	\]
    	Then for $k>k_\star$, $\rho_k\le C N^{-k_3}$ and we decompose
    	\[
    	\sum_{x\ne y}\rho_{xy}\,H(\rho_{xy})
    	\;=\;
    	\underbrace{\sum_{1\le k\le k_\star}
    		\sum_{\substack{\{x,y\}\subset V_N\\ d_{\G_N}(x,y)=k}}
    		\rho_k H(\rho_k)}_{=:\ \varepsilon_N^{\mathrm{near}}}
    	\;+\;
    	\underbrace{\sum_{k>k_\star}
    		\sum_{\substack{\{x,y\}\subset V_N\\ d_{\G_N}(x,y)=k}}
    		\rho_k H(\rho_k)}_{=:\ \varepsilon_N^{\mathrm{far}}}.
    	\]
    	For $1\le k\le k_\star$, using \eqref{eq:nk-rhok-bounds} and the monotonicity of $H$,
    	\[
    	\varepsilon_N^{\mathrm{near}}
    	\ \le\ \sum_{1\le k\le k_\star} n_k\,\rho_k\,H(\rho_k)
    	\ \le\ \sum_{1\le k\le k_\star} \big(C_1 N (r-1)^k\big)\big(C_2 (r-1)^{-k}\big)\,H(\rho_1)
    	\ \le\ C\,N\,k_\star\,H(\rho_1),
    	\]
    	where $\rho_1\le c_0<1$ depends only on $r$. 
    	As $a_N^2=2\log N+O(\log\log N)$, we find
    	\[
    	H(\rho_1)
    	=\frac{1}{\sqrt{1-\rho_1^2}}
    	\exp\!\Big(-\frac{a_N^2}{1+\rho_1}\Big)
    	\ \le\ C\,N^{-\,\frac{2}{1+\rho_1}}(\log N)^{O(1)}.
    	\]
    	Since $\frac{2}{1+\rho_1}>1$ for $r\ge3$, we conclude $\varepsilon_N^{\mathrm{near}}=o(1)$. For $k>k_\star$, $\rho_k\le C N^{-k_3}$, hence
    	\[
    	\varepsilon_N^{\mathrm{far}}
    	\ \le\ \binom{N}{2}\cdot 
    	\frac{C N^{-k_3}}{\sqrt{1-C^2N^{-2k_3}}}\,
    	\exp\!\Big(-\frac{a_N^2}{1+CN^{-k_3}}\Big)
    	\ =\ N^{-k_3+o(1)}\ =\ o(1).
    	\]
    	%
    	Combining both bounds, we see that $T_N(\G_N)=o(1)$, and thus by Theorem~\ref{thm:gauss-comp-ppp} we obtain
    	\[
    	\Big(P_{\widehat Z},\ \max_{x\in V_N}\tfrac{\widehat Z_x-a_N}{b_N}\Big)
    	\ \Longrightarrow\ (\mathcal P_\infty,\ \max\mathcal P_\infty).
    	\]
    	Recalling  $
    	$  that the terms $\delta_N=o(a_N^{-1})$ in \eqref{eq:witherror} are negligible, 
    	we obtain the same limit for the point processes
    	\(
        \sum_{x\in V_N}\delta_{(Z_x-a_N)/b_N}.
    	\)
    \end{proof}

\end{document}